# Sequential giant planet formation initiated by disc substructure

### v1

@@ 2.2 Planetesimal evolution
The realized planetesimals, and a Solar-mass star, are then evolved by SyMBAp with full gravitational interactions as well as additional subroutines to include gas drag, the planet-disc interactions, pebble accretion, gas accretion and planetary gap opening. If two bodies collide, they are assumed to merge completely, i.e., collisions are perfectly inelastic. At each communication step, on top of the newly formed planetesimals, the radial profiles of the disc are passed to SyMBAp. This includes the gas components:
@@
Also, the feedback to the disc is also passed to DustPy including
@@ 2.2.1 Pebble accretion
The treatment for the pebble accretion rate of each planetesimal is identical to that presented in Lau et al. (2022) and summarized below. First, the pebble mass flux of dust species $i$ at any location is given by
@@
We do not implement the pebble isolation mass explicitly with a prescription (e.g. Lambrechts et al., 2014) in this work. Instead, as dust and gas evolve consistently in this model, the planetary gap opening by a planet (Sect. 2.3) can interrupt the pebble flux capturing the process of pebble isolation within the model.
@@ 2.2.2 Gas accretion
We follow Piso & Youdin (2014) and Bitsch et al. (2015) to describe the gas accretion rate with the modification by Chambers (2021) to account for the energy released from pebble accretion. Gas accretion generally begins when the energy released from pebble accretion decreases enabling the cooling of the gas envelope. The gas accretion rate in this phase is
@@
with the opacity of the gas envelope $\kappa$ and the density of the core $\rho_{\mathrm{c}}=5.5\ \mathrm{g}\ \mathrm{cm}^{-3}$. We note that the solid mass accreted from planetesimal accretion is negligible compared to pebble accretion in our simulations. We follow Brouwers et al. (2021) for the grain size near the Bondi radius of the gas envelope to calculate its Rosseland mean opacity. Assuming the Epstein regime, the incoming solids converge to the size
@@ 2.2.3 Physical radius
As the planetesimals or planets grow by many orders of magnitude in mass, the physical radius $R_{\mathrm{p}}$ is evaluated correspondingly. For mass $m$ less than 0.1 Earth mass, we assume an internal density $\rho_{\mathrm{s}}$ of 1.5 $\mathrm{g}\ \mathrm{cm}^{-3}$, which is the same when the planetesimals are formed. For $m$ above 0.1 $M_{\oplus}$ but less than 5 $M_{\oplus}$, we follow Seager et al. (2007) for the mass-radius relationship of rocky planets, which is
@@
with the radius of Earth $R_{\oplus}$. For $m$ above 5 $M_{\oplus}$, we follow the mass-radius relationship applied in Matsumura et al. (2017), which is
@@ 2.3 Planetary gap opening
For all planetesimals and planets, the non-dimensional gap opening factor (Kanagawa et al., 2015) is evaluated, which is given by
@@
Since the treatment in Sect. 2.1.3 to attain a target gas profile does not change the actual turbulence experienced by the dust and we only consider the torque exerted on the gas disc, planetary gap opening is applied through the modified-$\alpha^{\prime}$ parameter. When $K>0.25$, we impose a planetary gap to the gas disc by dividing $\alpha^{\prime}$ by the ratio of the perturbed surface density to the unperturbed one $\Sigma_{\mathrm{g}}/\Sigma_{\mathrm{g},0}$. In other words, the planetary gap is imposed when the change caused by the corresponding body is more than about 1% of the unperturbed surface gas density.
The effects on $\alpha^{\prime}$ are multiplied when more than one planet can open a gap, i.e. for all gap opening planets $i$
$\alpha^{\prime}=\frac{\alpha}{F\cdot\prod_{i}\left(\Sigma_{\mathrm{g}}/\Sigma_{\mathrm{g},0}\right)_{i}}$ (26)
@@
We note that some small but short-period chaotic movements of massive planets can prevent the disc from converging to a quasi-steady state and the speed of code is significantly reduced. Therefore, we allow $\alpha^{\prime}(r)$ to reach the target value exponentially on a relaxation timescale of $250\times(10^{-3}/\alpha)$ years, which is below the viscous evolution timescales of the disc in the adopted radial domain of the simulations.
We adopt the empirical formula by Duffell (2020) for the gap profile, which is
@@ 2.4 Gas drag and planet-disc interactions
All bodies experience the combined effects of aerodynamic gas drag and the planet-disc interactions. For low-mass planet without gap opening, the treatment for gas drag and planet-disc interaction are identical to that presented in Lau et al. (2022) and summarized below.
@@
For type-I damping and migration, we adopt the prescription based on dynamical friction by Ida et al. (2020). The timescales for the isothermal case and finite $i$, while $i<\hat{h}_{\mathrm{g}}$, (Appendix D of Ida et al. (2020)) are implemented. The evolution timescales of semimajor axis, eccentricity and inclination are defined, respectively, by
@@
where $\Sigma_{\mathrm{g}}$ and $\hat{h}_{\mathrm{g}}$ are retrieved from the local radial cell of the disc model. The normalize torques $C_{\mathrm{M}}$ and $C_{\mathrm{T}}$ are given by
@@
with $p_{\Sigma}\equiv-\mathrm{d}\ln\Sigma_{\mathrm{g}}/\mathrm{d}\ln r$ and $q_{T}\equiv-\mathrm{d}\ln T/\mathrm{d}\ln r$. The three timescales are then applied to the equation of motion
@@
As the planet grows and opens a gap in the disc, Kanagawa et al. (2018) suggest that the magnitude of the torque scales linearly with the local surface density providing a smooth transition to the high-mass (type-II) regime of planet migration. Since the dependence of $t_{\mathrm{wav}}$ on $\Sigma_{\mathrm{g}}$ as in Eq. (35) is retrieved from the local grid cell, the above treatment combined with gap opening (Sect. 2.3) can already capture this transition.
We note that planetary gap opening (Sect. 2.3) and planet migration are the results of planet-disc interactions, which are physically coupled by the action-reaction pair. However, we have adopted two independent prescriptions for each of them since a prescription for a torque profile which is suitable for a 1-D model and a general planet mass is still missing. Further discussions on the adopted treatments are in Sect. 4.4.2.
@@ 2.5 Numerical setup
The time step for SyMBAp $\tau=0.2$ yr is used and particles are removed if the heliocentric distance is less than 4 au or greater than $100$ au. The additional subroutines for the evolution of the $N$-body particles are added to SyMBAp as
@@ 3.1 The case of $\alpha=5\times 10^{-4}$
Figure 3: Tracks of the massive bodies in the simulation with $\alpha=5\times 10^{-4}$. Each colour shows one of the five simulations corresponding to those in the final panel of Fig. 2, where the key timestamps are also denoted along the time axis on the right. The solid lines show the the semimajor axis of the bodies reached 10 $M_{\oplus}$ by the end of the simulations. The shaded areas indicate the extents of the lower and upper quartiles of the semimajor axes of all bodies when the total number is above 50.
@@ 3.1.1 Formation and evolution of massive bodies
At 0.34 Myr (iii), the first massive core has entered the runaway gas accretion phase and opened a significant gap in the disc as it becomes a gas giant ($>100M_{\oplus}$). The less massive core and planetesimals, which are also formed from the initial pressure bump, are being scattered mainly to wider orbits as shown by the tracks in Fig. 3. While most planetesimals have been scattered out of the system, the orbit of the second-most massive planetary core is circularized near 8 to 9 au and it continues to grow.
@@
Across the five random simulations, the final panel of Fig. 2 and Fig. 3 show very similar results for the gas giant-pair formed in the first generation. For the next generation of planet formation, further stochasticity presents. Two (blue and green) simulations produce a pair of ice giants and another two (red and purple) produce only one ice giant. One simulation (orange) shows no ice giant at 2 Myr but a swarm of still-growing planetesimals and planet embryos.
@@ 3.1.2 Dust mass budget
Figure 4: a) Solid mass budget in the simulation with $\alpha=5\times 10^{-4}$ that corresponds to the one presented by the colour blue in Fig. 2 & 3. The six timestamps in Fig. 2 are denoted on the time axis. The solid mass is divided into four categories: the dust mass $M_{\mathrm{d}}$ inside and outside of 5 au, solids bound in massive bodies and ejected massive bodies. It shows a high planet formation efficiency that the majority of solid mass (85% of the initial dust mass beyond 5 au) are converted into massive bodies. b) Cumulative inflow of dust entered 5 au after 0.5 Myr. A total of about 1.6 $M_{\oplus}$ of inflow to the inner disc is recorded over the next 1.5 Myr up to the end of the simulation.
@@
Shortly after 0.2 Myr, there is a spike in the dust mass inside of 5 au. This is caused by the gap opening in a dust rich location as the first planetary core reaches the mass of $10\ M_{\oplus}$. Similarly, another spike occurs at about 0.34 Myr (iii) when the first core enters the runaway gas accretion phase but it is much smaller as dust is already depleted around the planet. Pebble accretion resumes for the second core at about 0.4 Myr after its orbit has been circularized, which also causes the change in the dust mass.
@@ 3.2 The case of $\alpha=3\times 10^{-4}$
Figure 6 summarizes the final architecture of the planets where three out of the five simulations form three gas giants and the remaining two (blue and purple) form two. Also, a significant dust ring remains respectively for all simulations external to their outermost gas giant.
@@ 4.1 Sequential planet formation
Comparing the results with the ‘inside-out’ planet formation scenario by Chatterjee & Tan (2013) for Kepler systems, we note that gas giant that has reached its final mass is more likely to trigger the next generation of planet formation. For low-mass planets that cannot open a significant gap in the disc ($\lesssim 100M_{\oplus}$), dust leakage from the pressure bump at the outer edge of the gap is significant and requires a large supply from the outer disc to reach the conditions for planetesimal formation. Even if planetesimals may form, they are under greater gravitational influence of the planets as the width of the gap scales with $m^{1/2}$ (Kanagawa et al., 2016; Duffell, 2020) while the Hill radius scales with $m^{1/3}$. In this case, the perturbation from the planet is more likely to prevent the growth of the planetesimals as pebble accretion, particularly when the planet continues to grow, is sensitive to the relative velocity as noted in Lau et al. (2022). These planetesimals will likely be scattered out of the system as well, if the planet enters the runaway gas accretion phase to become a gas giant. Therefore, the outer edge of the planetary gap opened by a steady gas giant is a much more favourable environment for the next generation of planet formation.
@@ 4.3.1 Predictions of planet formation at pressure bumps
The case of fast core formation but slow gas accretion is suggested to form a chain of super-Earths or potentially Saturn-mass planets over a prolonged period of time. This case is similar to the scenario proposed by Chatterjee & Tan (2013). As discussed in Sect. 4.1, the planetary gap formed by super-Earths may not be able to trap a significant amount of dust and trigger planetesimal formation. And, even if planetesimals could form, they are likely much closer to the core and cannot grow efficiently by pebble accretion.
@@
Although dust trap favours planetesimal formation, we emphasize that in-situ planet formation at pressure bumps is unlikely a general solution to the diversity of exoplanetary systems. In particular, for the observed compact planet chains in resonance, a more probable formation scenario is that the cores are formed at a temporary pressure bump which migrate subsequently through the disc. Multiple works have studied the scenario that the inner most planet can be trapped at the disc edge (e.g. Terquem & Papaloizou, 2007; Cossou et al., 2013; Brasser et al., 2018; Huang & Ormel, 2023) and the subsequent inward-migrating planets can form a resonant chain through convergent inward migration (e.g. Tamayo et al., 2017; Delisle, 2017; MacDonald & Dawson, 2018; Wong & Lee, 2024).
@@ 4.3.2 Sandwiched planet formation
With gas and dust hydrodynamics simulations, Pritchard et al. (2024) proposed the ‘sandwiched planet formation’ scenario where planet formation can occur with the dust trapped between two massive planets that each creates a pressure maximum. The authors already noted that formation of the planets and dust fragmentation are not modelled, which may have critical effect to the dust concentration between the planets. Furthermore, from the results in the presented work, we also note that if planetesimals could form between the planets, they are likely under a great gravitational influence from the massive planets. This will prevent them from growing efficiently by accreting pebbles, or, more likely, scatter them. Nonetheless, their work confirms dust rings can be created by massive planets with hydrodynamics simulations and these are preferred locations of planet formation.
@@ 4.4.1 Initial disc substructure
In this work, we study the consequence of a substructure in the disc that can trigger planetesimal formation, with the location motivated by that of Jupiter. Although multiple non-planetary mechanisms are proposed in the field as discussed in Sect. 1, the parameter space, including the location, amplitude, width and lifetime, is not explored in this work and requires future investigations. For instance, from some test runs, we note that the amplitude of the substructure needs to be large enough to trap dust effectively and trigger planetesimal formation, although non-axisymmetric features that may aid dust concentration such as vortices (e.g. Barge & Sommeria, 1995; Tanga et al., 1996) are not considered. Sequential planet formation also cannot occur if the dust mass remaining is not enough to trigger the next generation of planet formation. We emphasize that the criteria to form planetesimals are not trivial to satisfy in typical disc conditions. This work only focuses on a case where planetesimal formation is possible. However, by combining the unknowns in the shape and location of the initial disc substructure with different disc parameters, we expect the model can produce a variety of planetary systems through a parameter study, where sequential planet formation may not always occur.

### v2

@@ 2.2 Planetesimal evolution
The realised planetesimals, and a Solar-mass star, were then evolved by SyMBAp with full gravitational interactions as well as additional subroutines to include gas drag, the planet-disc interactions, pebble accretion, gas accretion and planetary gap opening. If two bodies collide, they were assumed to merge completely, that is, collisions are perfectly inelastic. At each communication step, on top of the newly formed planetesimals, the radial profiles of the disc were passed to SyMBAp. This included the gas components:
@@
Also, the feedback to the disc was also passed to DustPy including
@@ 2.2.1 Pebble accretion
The treatment for the pebble accretion rate of each planetesimal is identical to that presented in Lau et al. (2022) and summarised below. First, the pebble mass flux of dust species $i$ at any location is given by
@@
We did not implement the pebble isolation mass explicitly with a prescription (e.g. Lambrechts et al., 2014) in this work. Instead, as dust and gas evolve consistently in this model, the planetary gap opening by a planet (Sect. 2.3) can interrupt the pebble flux capturing the process of pebble isolation within the model.
@@ 2.2.2 Gas accretion
We followed Piso & Youdin (2014) and Bitsch et al. (2015) to prescribe the gas accretion rate with the modification by Chambers (2021) to account for the energy released from pebble accretion. Gas accretion generally begins when the energy released from pebble accretion decreases enabling the cooling of the gas envelope. The gas accretion rate in this phase is
@@
with the opacity of the gas envelope $\kappa$ and the density of the core $\rho_{\mathrm{c}}=5.5\ \mathrm{g}\ \mathrm{cm}^{-3}$. We note that the solid mass accreted from planetesimal accretion is negligible compared to pebble accretion in our simulations. We followed Brouwers et al. (2021) for the grain size near the Bondi radius of the gas envelope to calculate its Rosseland mean opacity. Assuming the Epstein regime, the incoming solids converge to the size
@@ 2.2.3 Physical radius
As the planetesimals or planets grow by many orders of magnitude in mass, the physical radius $R_{\mathrm{p}}$ is evaluated correspondingly. For mass $m$ less than 0.1 Earth mass, we assumed an internal density $\rho_{\mathrm{s}}$ of 1.5 $\mathrm{g}\ \mathrm{cm}^{-3}$, which is the same when the planetesimals are formed. For $m$ above 0.1 $M_{\oplus}$ but less than 5 $M_{\oplus}$, we followed Seager et al. (2007) for the mass-radius relationship of rocky planets, which is
@@
with the radius of Earth $R_{\oplus}$. For $m$ above 5 $M_{\oplus}$, we followed the mass-radius relationship applied in Matsumura et al. (2017), which is
@@ 2.3 Planetary gap opening
For all planetesimals and planets, the non-dimensional gap opening factor (Kanagawa et al., 2015) was evaluated, which is given by
@@
Since the treatment in Sect. 2.1.3 to attain a target gas profile does not change the actual turbulence experienced by the dust and we only consider the torque exerted on the gas disc, planetary gap opening was applied through the modified $\alpha^{\prime}$-parameter. When $K>0.25$, we impose a planetary gap to the gas disc by dividing $\alpha^{\prime}$ by the ratio of the perturbed surface density to the unperturbed one $\Sigma_{\mathrm{g}}/\Sigma_{\mathrm{g},0}$. In other words, the planetary gap is imposed when the change caused by the corresponding body is more than about 1% of the unperturbed surface gas density.
The effects on $\alpha^{\prime}$ were multiplied when more than one planet can open a gap, that is, for all gap opening planets $i$
$\alpha^{\prime}=\frac{\alpha}{F\cdot\prod_{i}\left(\Sigma_{\mathrm{g}}/\Sigma_{\mathrm{g},0}\right)_{i}},$ (26)
@@
We note that some small but short-period chaotic movements of massive planets can prevent the disc from converging to a quasi-steady state and the speed of code is significantly reduced. Therefore, we allowed $\alpha^{\prime}(r)$ to reach the target value exponentially on a relaxation timescale of $250\times(10^{-3}/\alpha)$ years, which is below the viscous evolution timescales of the disc in the adopted radial domain of the simulations.
We adopted the empirical formula by Duffell (2020) for the gap profile, which is
@@ 2.4 Gas drag and planet-disc interactions
All bodies experience the combined effects of aerodynamic gas drag and the planet-disc interactions. For low-mass planet without gap opening, the treatment for gas drag and planet-disc interaction are identical to that presented in Lau et al. (2022) and summarised below.
@@
For type-I damping and migration, we adopted the prescription based on dynamical friction by Ida et al. (2020). The timescales for the isothermal case and finite $i$, while $i<\hat{h}_{\mathrm{g}}$, (Appendix D of Ida et al. (2020)) were implemented. The evolution timescales of semimajor axis, eccentricity and inclination are defined, respectively, by
@@
where $\Sigma_{\mathrm{g}}$ and $\hat{h}_{\mathrm{g}}$ are retrieved from the local radial cell of the disc model. The normalised torques $C_{\mathrm{M}}$ and $C_{\mathrm{T}}$ are given by
@@
with $p_{\Sigma}\equiv-\mathrm{d}\ln\Sigma_{\mathrm{g}}/\mathrm{d}\ln r$ and $q_{T}\equiv-\mathrm{d}\ln T/\mathrm{d}\ln r$. The three timescales were then applied to the equation of motion
@@
As the planet grows and opens a gap in the disc, Kanagawa et al. (2018) suggested that the magnitude of the torque scales linearly with the local surface density providing a smooth transition to the high-mass (type-II) regime of planet migration. Since the dependence of $t_{\mathrm{wav}}$ on $\Sigma_{\mathrm{g}}$ as in Eq. (35) is retrieved from the local grid cell, the above treatment combined with gap opening (Sect. 2.3) can already capture this transition.
We note that planetary gap opening (Sect. 2.3) and planet migration are the results of planet-disc interactions, which are physically coupled by the action-reaction pair. However, we have adopted two independent prescriptions for each of them since a prescription for a torque profile which is suitable for a one-dimensional model and a general planet mass is still missing. Further discussions on the adopted treatments are in Sect. 4.4.2.
@@ 2.5 Numerical setup
The time step for SyMBAp $\tau=0.2$ yr was used and particles were removed if the heliocentric distance is less than 4 au or greater than $100$ au. The additional subroutines for the evolution of the $N$-body particles were added to SyMBAp as
@@ 3.1 The case of $\alpha=5\times 10^{-4}$
Figure 3: Tracks of the massive bodies in the simulation with $\alpha=5\times 10^{-4}$. Each colour shows one of the five simulations corresponding to those in the final panel of Fig. 2, where the key timestamps are also denoted along the time axis on the right. The solid lines show the semimajor axis of the bodies reached 10 $M_{\oplus}$ by the end of the simulations. The shaded areas indicate the extents of the lower and upper quartiles of the semimajor axes of all bodies when the total number is above 50.
@@ 3.1.1 Formation and evolution of massive bodies
At 0.34 Myr (iii), the first massive core has entered the runaway gas accretion phase and opened a significant gap in the disc as it becomes a gas giant ($>100M_{\oplus}$). The less massive core and planetesimals, which are also formed from the initial pressure bump, are being scattered mainly to wider orbits as shown by the tracks in Fig. 3. While most planetesimals have been scattered out of the system, the orbit of the second-most massive planetary core is circularised near 8 to 9 au, and it continues to grow.
@@
Across the five random simulations, the final panel of Fig. 2 and Fig. 3 show very similar results for the gas giant pair formed in the first generation. For the next generation of planet formation, further stochasticity presents. Two (blue and green) simulations produce a pair of ice giants and another two (red and purple) produce only one ice giant. One simulation (orange) shows no ice giant at 2 Myr but a swarm of still-growing planetesimals and planet embryos.
@@ 3.1.2 Dust mass budget
Figure 4: Evolution of the solid mass in the simulation with $\alpha=5\times 10^{-4}$ that corresponds to the one presented by the colour blue in Fig. 2 & 3. The six timestamps in Fig. 2 are also denoted on the time axis. a) Solid mass budget. The solid mass is divided into four categories: the dust mass $M_{\mathrm{d}}$ inside and outside of 5 au, solids bound in massive bodies and ejected massive bodies. It shows a high planet formation efficiency that the majority of solid mass (85% of the initial dust mass beyond 5 au) are converted into massive bodies. b) Cumulative inflow of dust entered 5 au after 0.5 Myr. A total of about 1.6 $M_{\oplus}$ of inflow to the inner disc is recorded over the next 1.5 Myr up to the end of the simulation.
@@
Shortly after 0.2 Myr, there is a spike in the dust mass inside of 5 au. This is caused by the gap opening in a dust rich location as the first planetary core reaches the mass of $10\ M_{\oplus}$. Similarly, another spike occurs at about 0.34 Myr (iii) when the first core enters the runaway gas accretion phase but it is much smaller as dust is already depleted around the planet. Pebble accretion resumes for the second core at about 0.4 Myr after its orbit has been circularised, which also causes the change in the dust mass.
@@ 3.2 The case of $\alpha=3\times 10^{-4}$
Figure 6 summarises the final architecture of the planets where three out of the five simulations form three gas giants and the remaining two (blue and purple) form two. Also, a significant dust ring remains respectively for all simulations external to their outermost gas giant.
@@ 4.1 Sequential planet formation
Comparing the results with the ‘inside-out’ planet formation scenario by Chatterjee and Tan (2013) for Kepler systems, we note that gas giant that has reached its final mass is more likely to trigger the next generation of planet formation. For low-mass planets that cannot open a significant gap in the disc ($\lesssim 100M_{\oplus}$), dust leakage from the pressure bump at the outer edge of the gap is significant and requires a large supply from the outer disc to reach the conditions for planetesimal formation. Even if planetesimals may form, they are under greater gravitational influence of the planets as the width of the gap scales with $m^{1/2}$ (Kanagawa et al., 2016; Duffell, 2020) while the Hill radius scales with $m^{1/3}$. In this case, the perturbation from the planet is more likely to prevent the growth of the planetesimals as pebble accretion, particularly when the planet continues to grow, is sensitive to the relative velocity as noted in Lau et al. (2022). These planetesimals will likely be scattered out of the system as well, if the planet enters the runaway gas accretion phase to become a gas giant. Therefore, the outer edge of the planetary gap opened by a steady gas giant is a much more favourable environment for the next generation of planet formation.
@@ 4.3.1 Predictions of planet formation at pressure bumps
The case of fast core formation but slow gas accretion is suggested to form a chain of super-Earths or potentially Saturn-mass planets over a prolonged period of time. This case is similar to the scenario proposed by Chatterjee and Tan (2013). As discussed in Sect. 4.1, the planetary gap formed by super-Earths may not be able to trap a significant amount of dust and trigger planetesimal formation. And, even if planetesimals could form, they are likely much closer to the core and cannot grow efficiently by pebble accretion.
@@
Although dust trap favours planetesimal formation, we emphasise that insitu planet formation at pressure bumps is unlikely a general solution to the diversity of exoplanetary systems. In particular, for the observed compact planet chains in resonance, a more probable formation scenario is that the cores are formed at a temporary pressure bump which migrate subsequently through the disc. Multiple works have studied the scenario that the inner most planet can be trapped at the disc edge (e.g. Terquem & Papaloizou, 2007; Cossou et al., 2013; Brasser et al., 2018; Huang & Ormel, 2023) and the subsequent inward-migrating planets can form a resonant chain through convergent inward migration (e.g. Tamayo et al., 2017; Delisle, 2017; MacDonald & Dawson, 2018; Wong & Lee, 2024).
@@ 4.3.2 Sandwiched planet formation
With gas and dust hydrodynamics simulations, Pritchard et al. (2024) proposed the ‘sandwiched planet formation’ scenario where planet formation can occur with the dust trapped between two massive planets that each creates a pressure maximum. The authors already noted that formation of the planets and dust fragmentation are not modelled, which may have critical effect to the dust concentration between the planets. Furthermore, from the results presented above (Sect. 3), we also note that if planetesimals could form between the planets, they are likely under a great gravitational influence from the massive planets. This will prevent them from growing efficiently by accreting pebbles, or, more likely, scatter them. Nonetheless, their work confirms dust rings can be created by massive planets with hydrodynamics simulations and these are preferred locations of planet formation.
@@ 4.4.1 Initial disc substructure
In this work, we studied the consequence of a substructure in the disc that can trigger planetesimal formation, with the location motivated by that of Jupiter. Although multiple non-planetary mechanisms are proposed in the field as discussed in Sect. 1, the parameter space, including the location, amplitude, width and lifetime, is not explored in this work and requires future investigations. For instance, from some test runs, we note that the amplitude of the substructure needs to be large enough to trap dust effectively and trigger planetesimal formation, although non-axisymmetric features that may aid dust concentration such as vortices (e.g. Barge & Sommeria, 1995; Tanga et al., 1996) are not considered. Sequential planet formation also cannot occur if the dust mass remaining is not enough to trigger the next generation of planet formation. We emphasise that the criteria to form planetesimals are not trivial to satisfy in typical disc conditions. This work only focuses on a case where planetesimal formation is possible. However, by combining the unknowns in the shape and location of the initial disc substructure with different disc parameters, we expect the model can produce a variety of planetary systems through a parameter study, where sequential planet formation may not always occur.